OpenStudy (anonymous):

Need someone uber smart to have the patience and time to explain this to me :P Math Analysis: Prove the given identity tan x/1 + sec x + 1 + sec x/ tan x = 2 csc x

OpenStudy (rogue):

You need to have patience as well :P Don't keep spamming the same question, wait for someone to answer.

OpenStudy (rogue):

Is this what you have to prove?\[\tan x + \sec x + 1 + \frac {\sec x}{tanx} = 2 \csc x\]

OpenStudy (anonymous):

lol sorry I've been here since the afternoon getting homework help..

OpenStudy (anonymous):

NOO!! tan x/1+sec x + 1 + sec x/tan x = 2 csc x =p

OpenStudy (rogue):

Its hard to understand, can you use parenthesis to clarify please?

OpenStudy (anonymous):

(tan x/1) + sec x + 1 + (sec x/ tan x) = 2 csc x there :P

OpenStudy (rogue):

but isn't tan x/1 just tan x?

OpenStudy (anonymous):

no.

OpenStudy (anonymous):

im guessing ur not that good at math.. >w>

OpenStudy (anonymous):

tan x = sin x/cos x everyone knows this!!!

OpenStudy (rogue):

The sec x /tan x term is just csc x, so you have to show that the rest of the stuff simplifies to another csc x because their sums must be 2csc x. I can't tell what the other terms are, so I can't help much with that, sorry.

OpenStudy (anonymous):

that's ok i knew you couldnt help me in the first place :P

OpenStudy (rogue):

.... You little..........

OpenStudy (rogue):

Well, have fun with it ;D